OpenStudy (anonymous):

an engineer on the ground is looking at the top of a building. the angle of elevation to the top of the building is 46 degrees... the engineer know the building is 250 ft tall... what is the distance from the engineer to the base of the building to the nearest whole foot?

OpenStudy (amistre64):

tan(46) = 250/d d = 250/tan(46)